# FusionRF: High-Fidelity Satellite Neural Radiance Fields from Multispectral and Panchromatic Acquisitions

### v2

@@ Abstract
We introduce FusionRF, a novel framework for digital surface reconstruction from satellite multispectral and panchromatic images. Current work has demonstrated the increased accuracy of neural photogrammetry for surface reconstruction from optical satellite images compared to algorithmic methods. Common satellites produce both a panchromatic and multispectral image, which contain high spatial and spectral information respectively. Current neural reconstruction methods require multispectral images to be upsampled with a pansharpening method using the spatial data in the panchromatic image. However, these methods may introduce biases and hallucinations due to domain gaps. FusionRF introduces joint image fusion during optimization through a novel cross-resolution kernel that learns to resolve spatial resolution loss present in multispectral images. As input, FusionRF accepts the original multispectral and panchromatic data, eliminating the need for image preprocessing. FusionRF also leverages multimodal appearance embeddings that encode the image characteristics of each modality and view within a uniform representation. By optimizing on both modalities, FusionRF learns to fuse image modalities while performing reconstruction tasks and eliminates the need for a pansharpening preprocessing step. We evaluate our method on multispectral and panchromatic satellite images from the WorldView-3 satellite in various locations, and show that FusionRF provides an average of 17% improvement in depth reconstruction accuracy, and renders sharp training and novel views.
Index Terms:
Neural Rendering, Photogrammetry, Satellite Imagery, Image Fusion, Pansharpening, Digital Surface Modeling, Remote Sensing, Deep Learning
@@ I Introduction
DIGITAL Surface Models (DSMs) allow us to visualize and understand the topology of our world. Historically, elevation data used to construct DSMs has been collected with complex and expensive specialized sensors and satellites, such as SAR [13] and LiDAR [14]. Recent work has attempted to reconstruct accurate surface models and DSMs from much more commonly available optical satellite imagery data, collected from various angles and over the span of years or months. By generating accurate surface reconstructions, these methods can visualize the scene from novel view angles even under different weather conditions.
@@
While algorithmic methods [15, 16] depend on calibrated stereo or tri-stereo image products, neural rendering methods [20, 21, 22, 23] instead leverage larger datasets of commercially available satellite images. These methods, known as Neural Radiance Fields (NeRF), perform scene reconstruction and surface modeling by optimizing multilayer perceptrons (MLPs) to represent the density and radiance information of a scene, creating an implicit understanding of the scene geometry. The MLPs can then be sampled from various input views for inference, generating depth maps and novel views of the scene. NeRFs have been shown to outperform previous algorithmic methods in surface reconstruction while introducing the ability to recognize and remove transient objects [21], model the scene under various sun angles [20], and estimate bundle adjustment parameters [23]. Additionally, NeRFs have the ability to model photometric differences between views [18] and resolve image quality issues such as camera blur and loss of resolution [19, 33, 40].
@@
Current NeRF architectures for satellite reconstruction [22, 23] depend on pansharpening prior to optimization to create ground truth imagery. Pansharpening methods can produce variances in upsampled images and often lack generalizability across satellite image sensors and unseen domains [24]. Since a satellite neural reconstruction model uses the upsampled images as ground truth during optimization, any hallucinations will also be present in the final reconstruction and affect the accuracy of the Digital Surface Model.
@@
Our proposed method, FusionRF, is a neural radiance field which optimizes directly on the full-channel multispectral and panchromatic images from satellite acquisitions. During optimization, fusion is facilitated by a cross-resolution kernel that simulates the loss of resolution in multispectral images, while a modal appearance embedding encodes the characteristics of panchromatic and multispectral views. Image appearance embeddings are employed to resolve photometric differences between images, while transient embeddings in conjunction with uncertainty learning help separate transient phenomena such as cars, construction activity, and trees from static objects in the scene. FusionRF learns to perform image fusion during training, producing high-fidelity novel view synthesized images. During evaluation, the cross-resolution kernel is disabled, allowing the NeRF to produce high-resolution novel views which are consistent with the input scene and outperform publicly available state-of-the-art methods in 3D reconstruction tasks.
@@ II-A Pansharpening
Commercial satellites are typically equipped with a pair of sensors that separately gather detailed spatial and spectral data. Some applications, such as water resource management [9] and land use management [1, 2, 4] require gathered light to be split into narrow frequency bands to analyze and monitor water quality and vegetation health. Multispectral sensors filter incoming light into these specialized channels and require larger photosites, trading spacial fidelity for spectral resolution. Panchromatic sensors accept one broad band of light, creating a single band image with high spatial resolution.
@@ II-C Resolution Recovery NeRF Variants
[…]
@@ III Method
Figure 2: Our Network Architecture: Every input ray $r$ is projected from origin o in $I_{ms}$ and $I_{pan}$ to the same ground point $g$ The sparse kernel $F_{\Phi}$ predicts weights $w_{q}$ for static locations $\{q\}$ surrounding o, which are then combined with the color predictions $\{c_{q}\}$ to produce the final output color.
@@ III-A RPC-based Sampling
Similarly to Sat-NeRF [21], the origin o of each ray is calculated by georeferencing the origin pixel $(x,y)$ with $h_{max}$. The destination g of each ray can be calculated in kind by georeferencing $(x,y)$ with $h_{min}$. The origin o and destination g are converted to a geocentric coordinate system, and the direction d of each ray is calculated as:
$\textbf{d}=\frac{\textbf{g}-\textbf{o}}{||\textbf{g}-\textbf{o}||^{2}}.$ (3)
The ray parameters o and d are then normalized to reduce the required precision for the model. Each ray can then be sampled between 0 and $||\textbf{g}-\textbf{o}||^{2}$.
@@ III-B Cross-Resolution Kernel
We introduce a novel cross-resolution kernel $F_{\Psi}$ to model the resolution difference between image $I_{ms}$ and image $I_{pan}$. Optimizing a NeRF over $I_{ms}$ would require that the model predict the same low resolution $I_{ms}$ for each ray originating in the image. As we wish to prevent our model from learning directly from $I_{ms}$, we estimate a relationship $I_{ms}=(I_{pred}\circledast F_{\Psi})$ where $I_{pred}$ represents the image predicted by the NeRF $F_{\theta}$. Since the datasets for satellite reconstruction tasks contain relatively few images, we design a kernel $F_{\Psi}$ with minimal parameters and a fixed sparse support of 9 rays in a $12r{+}1\times 2r{+}$ 1 window.
For a pixel $p$, scaling factor $r$, and coordinates $(x,y)$, we define the fixed offsets:
@@
Empirically, we find that the inclusion of $F_{\Psi}$ in this fashion allows the multispectral images rendered without a sparse kernel to be considerably sharper. A more complex design of $F_{\Psi}$, such as randomizing the positions or predicting deformations of q, leads to an over-parameterized model and negatively impacts reconstruction accuracy.
@@ III-C Intrinsic Pansharpening
Figure 3: Embedding Diagram: One modal embedding is shared across all panchromatic and multispectral each, allowing the model to encode modal information in a uniform representation. Each image view, represented by date of capture, also shares one image embedding across both the panchromatic and multispectral images. In the case of a view containing only one modality, the image embedding is unique to that image.
The images $I_{lms}$ and $I_{pan}$ are collected simultaneously and aligned during collection. Therefore, a pixel $p$ at location $(x,y)$ describes the same geographic location g in both images. FusionRF takes both images as input independently of each other, creating two rays $\textbf{r}_{ms}$ and $\textbf{r}_{pan}$ which both begin at $(x,y)$ and describe g. These rays differ in their predicted colors $\hat{\textbf{C}}(\textbf{r})$: $\textbf{r}_{ms}$ predicts the color of $p$ in $I_{lms}$ and $\textbf{r}_{pan}$ in $I_{pan}$. Therefore, the loss computation for each image will be independent of each other, but will each be calculated against the same g. For the multispectral image, the loss $\mathcal{L}_{Image}=||\hat{\textbf{C}}({\textbf{r}})-\textbf{C}_{lms}({\textbf{r}})||^{2}$ encourages the model to predict a $\hat{\textbf{C}}(\textbf{r})$ with a blurry color and good multispectral resolution across the full 8 output channels. In contrast, the loss for the panchromatic image $\mathcal{L}_{Image}=||\hat{\textbf{C}}({\textbf{r}})-\textbf{C}_{pan}({\textbf{r}})||^{2}$ encourages the model to predict a sharp image with panchromatic color. Here, $\textbf{C}_{lms}({\textbf{r}})$ and $\textbf{C}_{pan}({\textbf{r}})$ represent the ground truth color for the pixel $p$ in the input images $I_{lms}$ and $I_{pan}$.
@@
It is important to note that there is no requirement on the availability of panchromatic data. While a traditional pansharpening method depends on the panchromatic pair $I_{pan}$ for a multispectral image $I_{ms}$ to perform sharpening, our model is able to share information between images from the same scene through the $F_{\Theta}$ MLP to fill in for a missing $I_{pan}$ pair. Information from the other available $I_{pan}$ images in the scene can still be applied to improve the sharpness of the rendered view of $I_{ms}$. Therefore, intrinsic pansharpening within the training dataset is still possible with limited panchromatic data and can even be done on novel multispectral views of the scene.
Our approach, shown graphically in Figure 2, encourages the model to develop images that exactly match the original $I_{ms}$ and $I_{pan}$, removing the requirements for hand-crafted image fusion metrics or artificial downsampling of training data. Rather than a separate pre-processing operation, pansharpening becomes an internal task for our model to solve in the production of high quality novel views. The model’s primary task is novel view synthesis, which ensures the predictions are grounded in the observed scene, reducing the likelihood of hallucinations. Since only one scene is considered in the training stage, the model cannot carry over artifacts from the domain of other scenes.
@@ III-D Multispectral NeRF
To encourage $F_{\Theta}$ to disambiguate between panchromatic and multispectral images in our dataset, we incorporate a modal embedding $t_{m}$ of size $(2,n)$. For each ray, $t_{m}$ is indexed to return the specific length $n$ embedding for the ray’s modality.
@@
TABLE I: Pansharpening Results Comparison: All methods were tested on full-size input multispectral and panchromatic data. Best results bolded, second best underlined.
@@
R-Q2n $\uparrow$ R-ERGAS $\downarrow$ R-SAM $\downarrow$ Depth MAE $\downarrow$ R-Q2n $\uparrow$ R-ERGAS $\downarrow$ R-SAM $\downarrow$ Depth MAE $\downarrow$ R-Q2n $\uparrow$ R-ERGAS $\downarrow$ R-SAM $\downarrow$ Depth MAE $\downarrow$ R-Q2n $\uparrow$ R-ERGAS $\downarrow$ R-SAM $\downarrow$ Depth MAE $\downarrow$
@@
Figure 4: Pansharpening Visual Comparison: The first row shows the original multispectral and panchromatic images alongside FusionRF’s generated image and pansharpened images from deep learning methods. The images illustrate that pansharpening methods tend to hallucinate color casts and extra details on buildings and the variance in images produced by pansharpening methods. The second row shows the LIDAR depth map provided by the dataset, along with the rendered depth from a FusionRF model trained on the raw panchromatic and multispectral data along with FusionRF models trained on pansharpened imagery. This comparison shows that achieving optimal image quality does not improve depth reconstruction.
III-E Transient Embedding
Due to the nature of satellite orbits around the Earth, the collection times for a particular scene can span a long period. While appearance embedding attempts to capture the variations in overall lighting and photometry, a transient embedding is also employed to capture the uncertainty of each pixel in the scene [18, 21]. This is particularly useful in our case, as adapting the reconstruction loss to ignore areas of constant change such as parking lots, construction zones, and treetops allows the model to better refine static areas such as buildings. In our model the uncertainty $\beta$ is predicted as a function of previous layer features $h$ and transient embedding $t_{u}$ through layer $F_{\Theta}^{\text{Transient}}(\textbf{h},t_{u})=\beta$, similarly to NeRF-W [18] and SatNeRF [21]. The uncertainty $\beta$ attenuates the importance of a pixel in the network’s loss function:
$\mathcal{L}=\frac{||\hat{\textbf{C}}({\textbf{r}})-\textbf{C}_{lms}({\textbf{r}})||^{2}}{2\beta_{r}^{2}}+\frac{\log(\beta(\textbf{r})+\beta_{min})+\beta_{offset}}{2}.$ (10)
While the transient mask allows the model to selectively ignore pixels it believes belong to transient objects, it can also be abused to ignore areas that are simply difficult to reconstruct. As a safeguard, the balancing term $\log(\beta+\beta_{min})+\beta_{offset}$ is introduced to penalize the model excessively labeling uncertain regions to drive down the training loss.
IV Experiments
IV-A Pansharpening Effectiveness Comparison
We perform two experiments to comparatively evaluate the ability of FusionRF to replace the pansharpening pre-processing stage. First, we measure the depth reconstruction quality under two conditions: FusionRF (with no cross-resolution kernel $F_{\Psi}$) trained on pansharpened images from state-of-the-art deep learning methods, and FusionRF (with $F_{\Psi}$) trained on the original multispectral and panchromatic images. The pansharpening images were created by four state-of-the-art methods: DRPNN [53], FusionNet [27], ZS-Pan [69], and PS-Dip [70].
@@ IV-A Pansharpening Effectiveness Comparison
Columns R-Q2n, R-ERGAS, and R-SAM in Table I show that pansharpening algorithms achieve strong image fusion quality in human perceptual visual quality metrics. However, this performance does not correlate to an improvement in column Depth MAE which represents 3D reconstruction accuracy results as described in Section IV-B. The depth reconstruction performance for each dataset varies greatly depending on which pansharpening method is used. FusionRF trained on panchromatic and multispectral images with $F_{\Psi}$ enabled provides the best reconstruction performance.

### v3

@@ Abstract
We introduce FusionRF, a novel framework for digital surface reconstruction from satellite multispectral and panchromatic images. Current work has demonstrated the increased accuracy of neural photogrammetry for surface reconstruction from optical satellite images compared to algorithmic methods. Common satellites produce both a panchromatic and multispectral image, which contain high spatial and spectral information respectively. Current neural reconstruction methods require multispectral images to be upsampled with a pansharpening method using the spatial data in the panchromatic image. However, these methods may introduce biases and hallucinations due to domain gaps. FusionRF introduces joint image fusion during optimization through a novel cross-resolution kernel that learns to resolve spatial resolution loss present in multispectral images. As input, FusionRF accepts the original multispectral and panchromatic data, eliminating the need for image preprocessing. FusionRF also leverages multimodal appearance embeddings that encode the image characteristics of each modality and view within a uniform representation. By optimizing on both modalities, FusionRF learns to fuse image modalities while performing reconstruction tasks and eliminates the need for a pansharpening preprocessing step. We evaluate our method on multispectral and panchromatic satellite images from the WorldView-3 satellite in various locations, and show that FusionRF provides an average of 17% reduction in depth reconstruction error, and renders sharp training and novel views.
@@ I Introduction
DIGITAL Surface Models (DSMs) allow us to visualize and understand the topology of our world. Historically, elevation data used to construct DSMs has been collected with complex and expensive specialized sensors and satellites such as SAR [5] and LiDAR [6]. Recent work has attempted to reconstruct accurate surface models and DSMs from much more commonly available optical satellite imagery, collected from various angles and over the span of years or months. By generating accurate surface reconstructions, these methods can visualize the scene from novel view angles even under different weather conditions.
While algorithmic methods [7, 8] depend on calibrated stereo or tri-stereo image products, neural rendering methods [12, 13, 14, 15] leverage larger datasets of commercially available satellite images. These methods, known as Neural Radiance Fields (NeRF), perform scene reconstruction and surface modeling by optimizing multilayer perceptrons (MLPs) to represent the density and radiance information of a scene, creating an implicit understanding of the scene geometry. The MLPs can then be sampled from various input views for inference, generating depth maps and novel views of the scene. NeRFs have been shown to outperform previous algorithmic methods in surface reconstruction while introducing the ability to recognize and remove transient objects [13], model the scene under various sun angles [12], and estimate bundle adjustment parameters [15]. Additionally, NeRFs have the ability to model photometric differences between views [10] and resolve image quality issues such as camera blur and loss of resolution [11, 24, 26].
@@
Current NeRF architectures for satellite reconstruction [14, 15] depend on pansharpening prior to optimization to create training imagery. Pansharpening methods can produce variance in upsampled images and often lack generalizability across satellite image sensors and unseen domains [16]. Since a satellite neural reconstruction model uses the upsampled images as ground truth during optimization, any hallucinations will also be present in the final reconstruction and affect the accuracy of the Digital Surface Model.
@@
Our proposed method, FusionRF, is a neural radiance field which optimizes directly on the full-channel multispectral and panchromatic images from satellite acquisitions. During optimization, fusion is facilitated by a cross-resolution kernel that simulates the loss of resolution in multispectral images, while a modal appearance embedding encodes the characteristics of panchromatic and multispectral views. Image appearance embeddings are employed to resolve photometric differences between images, while transient embeddings in conjunction with uncertainty learning help separate transient phenomena such as cars, construction activity, and seasonal foliage from static objects in the scene. FusionRF learns to perform image fusion during training, producing high-fidelity novel view synthesized images. During evaluation, the cross-resolution kernel is disabled, allowing the NeRF to produce high-resolution novel views which are consistent with the input scene and outperform publicly available state-of-the-art methods in 3D reconstruction tasks.
@@ II-A Pansharpening
Commercial satellites are typically equipped with a pair of sensors that separately gather detailed spatial and spectral data. Some applications, such as water resource management [4] and land use management [1, 2, 3] require gathered light to be split into narrow frequency bands to analyze and monitor water quality and vegetation health. Multispectral sensors filter incoming light into these specialized channels and require larger photosites, trading spatial fidelity for spectral resolution. Panchromatic sensors accept one broad band of light, creating a single band image with high spatial resolution.
@@ II-B Neural Radiance Fields and Variants
Here, $T_{i}$ represents the transmittance, or the probability that the ray travels without interference, while $\delta_{i}$ represents the distance between samples. The predicted color $\hat{\textbf{C}}(\textbf{r})$ is weighted by the transmittance and opacity of the scene. This allows the color to be controlled by the predicted scene content along the entire ray, taking into account occlusions and reflections. The loss $\mathcal{L}$ is computed as the error between the colors generated by the rendered rays and the ground truth pixel color $\textbf{C}(\textbf{r})$:
@@ II-C Resolution Recovery NeRF Variants
Recent methods [11, 24, 26] address the problem of deblurring input images by modeling a blurry image as a convolution of a blur kernel with a sharp image. For a particular pixel $p$ with color $\textbf{c}_{p}$, the blurry color can be calculated as a convolution $\textbf{b}_{p}=\textbf{c}_{p}\circledast\textbf{H}$. The blur kernel H typically consists of a $K\times K$ matrix. However, while a typical deblurring model convolves a predicted image with H to compare to the blurry ground truth, a NeRF model cannot due to the computational complexity of rendering $K^{2}\times h\times w$ rays for an image with dimensions $(h,w)$. These methods seek to resolve this issue by approximating the dense blur kernel H. A small set of points are selected and rays $\{\textbf{q}\}$ are generated as a sparse approximation. A separate MLP then predicts the weights $w_{\textbf{q}}$, as well as regresses the optimal positions of the origins and destinations of q. Each ray q generates a color $\textbf{c}_{q}$ which is then weighted by $w_{\textbf{q}}$ to produce the blurry color: $\textbf{b}_{p}=\sum_{\textbf{q}}\textbf{c}_{\textbf{q}}w_{\textbf{q}}$. The deformation of rays allows for the kernel to model complex motion and defocus blur. Because the underlying NeRF model predicts sharp images which are then deformed by the blur kernel to match the blurry inputs, we can remove the kernel in evaluation to render sharp images of the scene.
@@ III Method
Figure 2: Our Network Architecture: Every input ray $r$ is projected from origin o in $I_{lms}$ and $I_{pan}$ to the same ground point $g$. The sparse cross-resolution kernel $F_{\Psi}$ predicts weights $w_{q}$ for static locations $\{q\}$ surrounding o, which are then combined with the color predictions $\{c_{q}\}$ to produce the final output color.
@@ III-A RPC-based Sampling
Similarly to Sat-NeRF [13], the origin o of each ray is calculated by georeferencing the origin pixel $(x,y)$ with $h_{max}$. The destination g of each ray can be calculated similarly by georeferencing $(x,y)$ with $h_{min}$. The origin o and destination g are converted to a geocentric coordinate system, and the direction d of each ray is calculated as:
[…]
The ray parameters o and d are then normalized to reduce the required precision for the model. Each ray can then be sampled between 0 and $\textbf{g}-\textbf{o}||^{2}$.
@@ III-B Cross-Resolution Kernel
We introduce a novel cross-resolution kernel $F_{\Psi}$ to model the resolution difference between image $I_{lms}$ and image $I_{pan}$. Optimizing a NeRF over $I_{lms}$ would require that the model predict the same low resolution $I_{lms}$ for each ray originating in the image. As we wish to prevent our model from learning directly from $I_{lms}$, we estimate a relationship $I_{lms}=(I_{pred}\circledast F_{\Psi})$ where $I_{pred}$ represents the image predicted by the NeRF $F_{\Theta}$. Since the datasets for satellite reconstruction tasks contain relatively few images, we design a kernel $F_{\Psi}$ with minimal parameters and a fixed sparse support of nine rays in a $2r{+}1\times 2r{+}1$ window. This pattern was chosen to align with the upscaling factor $r$ of the $I_{lms}$ image, placing each support ray $r$ pixels away in each dimension from all other support rays. This sampling strategy ensures that support rays avoid redundant information from neighboring pixels in $I_{lms}$ caused by upscaling, while capturing information from neighboring pixels in the original $I_{ms}$ image.
For a pixel $p$, scaling factor $r$, and coordinates $(x,y)$, we define the fixed positions:
@@
Empirically, we find that the inclusion of $F_{\Psi}$ in this fashion allows the multispectral images rendered without a sparse cross-resolution kernel to be considerably sharper. A more complex design of $F_{\Psi}$, such as randomizing the positions or predicting deformations of q, leads to an over-parameterized model and negatively impacts reconstruction accuracy.
@@ III-C Intrinsic Pansharpening
Figure 3: Embedding Diagram: One modal embedding is shared across all panchromatic images and another across all multispectral images, allowing the model to encode modal information in a uniform representation. Each image view, represented by date of capture, also shares one image embedding across both the panchromatic and multispectral images. In the case of a view containing only one modality, the image embedding is unique to that image.
[…]
@@
It is important to note that FusionRF can still operate with limited panchromatic imagery. While a traditional pansharpening method depends on the panchromatic pair $I_{pan}$ for a multispectral image $I_{lms}$ to perform sharpening, our model is able to share information between images from the same scene through the $F_{\Theta}$ MLP to fill in for a missing $I_{pan}$ pair. Information from the other available $I_{pan}$ images in the scene can still be applied to improve the sharpness of the rendered view of $I_{lms}$. Therefore, intrinsic pansharpening within the training dataset is still possible with limited panchromatic data and can even be done on novel multispectral views of the scene.
Our approach, shown graphically in Figure 2, encourages the model to develop images that exactly match the original $I_{lms}$ and $I_{pan}$, removing the requirements for hand-crafted image fusion metrics or artificial downsampling of training data. Rather than a separate pre-processing operation, pansharpening becomes an internal task for our model to solve in the production of high quality novel views. The model’s primary task is novel view synthesis, which ensures the predictions are grounded in the observed scene, reducing the likelihood of hallucinations. Since only one scene is considered in the training stage, the model cannot carry over artifacts from the domain of other scenes.
@@ III-D Multispectral NeRF
To encourage $F_{\Theta}$ to disambiguate between panchromatic and multispectral images in our dataset, we incorporate a modal embedding $t_{m}$ of size $(2,n)$. For each ray, $t_{m}$ is indexed to return the specific embedding of size $n$ for the ray’s modality.
@@
Figure 4: Pansharpening Visual Comparison: The first row shows the original multispectral and panchromatic images alongside FusionRF’s generated image and pansharpened images from deep learning methods. The images show the tendency of pansharpening methods to hallucinate color casts and additional details on buildings and shows the variance in the final result produced by pansharpening methods. The second row shows the LIDAR depth map provided by the dataset, along with the rendered depth from a FusionRF model trained on the raw panchromatic and multispectral data along with FusionRF models trained on pansharpened imagery. Finally, the third row shows the error maps of these reconstructions against the LIDAR depth map. This comparison shows that achieving optimal image quality does not improve depth reconstruction.
III-E Transient Embedding
Due to the nature of satellite orbits around the Earth, the collection times for a particular scene can span a long period. While appearance embedding attempts to capture the variations in overall lighting and photometry, a transient embedding is also employed to capture the uncertainty of each pixel in the scene [10, 13]. This is particularly useful in our case, as adapting the reconstruction loss to ignore areas of constant change such as parking lots, construction zones, and treetops allows the model to better refine static areas such as buildings. In our model the uncertainty $\beta$ is predicted as a function of previous layer features $h$ and transient embedding $t_{u}$ through layer $F_{\Theta}^{\text{Transient}}(\textbf{h},t_{u})=\beta$, similarly to NeRF-W [10] and Sat-NeRF [13]. The uncertainty $\beta$ attenuates the importance of a pixel in the network’s loss function:
[…]
While the transient mask allows the model to selectively ignore pixels it believes belong to transient objects, it can also be abused to ignore areas that are simply difficult to reconstruct. As a safeguard, the balancing term $\log(\beta+\beta_{min})+\beta_{offset}$ is introduced to penalize excessive labeling of uncertain regions to drive down the training loss.
Figure 5: Spectral Comparison: For each of the eight bands of the WorldView-3 image shown in Figure 4, we plot the spectral density in the form of a histogram of pixel values. We display results for the FusionRF generated views from the multispectral and panchromatic data as well as the images pansharpened by PSDip and DRPNN.
IV Experiments
IV-A Pansharpening Effectiveness Comparison
We perform two experiments to comparatively evaluate the ability of FusionRF to replace the pansharpening pre-processing stage. First, we measure the depth reconstruction quality under two conditions: FusionRF (with no cross-resolution kernel $F_{\Psi}$) trained on pansharpened images from state-of-the-art deep learning methods, and FusionRF (with $F_{\Psi}$) trained on the original multispectral and panchromatic images. The pansharpening images were created by four state-of-the-art methods: DRPNN [39], FusionNet [19], ZS-Pan [47], and PSDip [48].
TABLE I: Pansharpening Results Comparison: All methods were tested on full-size multispectral and panchromatic data. Best results bolded, second best underlined. Higher values are better for R-Q2n, and lower are better for R-ERGAS, R-SAM, and MAE.
R-Q2n : $\uparrow$       R-ERGAS, R-SAM, Depth MAE: $\downarrow$
@@ IV-A Pansharpening Effectiveness Comparison
R-Q2n R-ERGAS R-SAM Depth MAE R-Q2n R-ERGAS R-SAM Depth MAE R-Q2n R-ERGAS R-SAM Depth MAE R-Q2n R-ERGAS R-SAM Depth MAE
@@
Columns R-Q2n, R-ERGAS, and R-SAM in Table I show that pansharpening algorithms achieve strong image fusion quality in human perceptual visual quality metrics. However, this performance does not correlate to an improvement in column Depth MAE which represents 3D reconstruction accuracy results as described in Section IV-B. These results show a disconnect between optimization of images for human perception and depth reconstruction. The improvements in image quality index performance that pansharpening algorithms achieve do not necessarily result in an improvement in depth estimation performance. FusionRF trained on panchromatic and multispectral images with $F_{\Psi}$ enabled avoids this pansharpening algorithm selection problem by intrinsically pansharpening the input data, providing the best depth reconstruction performance while still being competitive in pansharpening metrics. Figure 5 shows the difference in spectral density between the two best performing pansharpening methods and the FusionRF result.
These results show that the chosen pansharpening algorithm significantly influences the performance of the downstream reconstruction algorithm, and intrinsic pansharpening through the $F_{\Psi}$ kernel provides the best reconstruction accuracy.